# Circuit complexity and functionality: a statistical thermodynamics perspective

### v2

@@
Circuit complexity and functionality:
a statistical thermodynamics perspective
Claudio Chamon Boston University    Andrei E. Ruckenstein    Eduardo R. Mucciolo University of Central Florida    Ran Canetti
@@
Circuit complexity, defined as the minimum circuit size required for implementing a particular Boolean computation, is a foundational concept in computer science. Determining circuit complexity is believed to be a hard computational problem [1]. Recently, in the context of black holes, circuit complexity has been promoted to a physical property, wherein the growth of complexity is reflected in the time evolution of the Einstein-Rosen bridge (“wormhole”) connecting the two sides of an AdS “eternal” black hole [2]. Here we explore another link between complexity and thermodynamics for circuits of given functionality, making the physics-inspired approach relevant to real computational problems, for which functionality is the key element of interest. In particular, our thermodynamic framework provides a new perspective on the obfuscation of programs of arbitrary length – an important problem in cryptography – as thermalization through recursive mixing of neighboring sections of a circuit, which can be viewed as the mixing of two containers with “gases of gates”. This recursive process equilibrates the average complexity and leads to the saturation of the circuit entropy, while preserving functionality of the overall circuit. The thermodynamic arguments hinge on ergodicity in the space of circuits which we conjecture is limited to disconnected ergodic sectors due to fragmentation. The notion of fragmentation has important implications for the problem of circuit obfuscation as it implies that there are circuits with same size and functionality that cannot be connected via local moves. Furthermore, we argue that fragmentation is unavoidable unless the complexity classes NP and coNP coincide, a statement that implies the collapse of the polynomial hierarchy of computational complexity theory to its first level.
Introduction
During the past decade a novel connection between physics and computer science has emerged in the course of explorations of one of the most fundamental open problems in physics, namely, the development of a theory of quantum gravity that reconciles and unifies general relativity with quantum mechanics. In this context, a bold proposal was made and developed by Susskind and collaborators [2, 3, 4] whereby the formal computer science notion of complexity has acquired specific physical reality, as it is conjectured that the growth of computational complexity represents the growth of the Einstein-Rosen bridge (the “wormhole”) of an AdS “eternal” black hole. It is argued that this growth is linear in time and persists for a time scale that is exponential in the physical black hole entropy. One way to formalize this connection is to use heuristic models of black hole dynamics based on quantum circuits of $2$-qubit gates [5], in which case one (i) defines “circuit complexity” as the minimum number of elementary gates needed to implement a particular unitary operator; and (ii) introduces an intuitive notion of “circuit entropy”, determined from the logarithm of the number of possible circuits of a given complexity. Unlike the physical entropy, which saturates to a value linear in the number of qubits, $n$, the circuit complexity, grows to its saturation value of $O(4^{n})$. More precisely, Refs. [2, 3, 4] argue that computational complexity is connected with the entropy of an auxiliary ensemble of $2^{n}$ classical particles moving on a negatively curved two-dimensional surface of large genus. In analogy with the second law of thermodynamics in physical systems, this connection leads to the natural description of the growth tendency of computational complexity as “the second law of complexity” [4]. It is important to note that in deriving the linear growth of complexity with number of gates, Refs. [3, 4] ignore the role of “circuit collisions”, defined as different sequences of gates that implement the same unitary transformation, i.e., circuits of the same “functionality.”  [^1] To eliminate any possible confusion with the notion of “gate collisions,” which we introduce later in the paper, “circuit collisions” in the sense of Refs. [3, 4] will be referred to as functionality degeneracies, which are quantified by the notion of circuit entropy defined below.
@@ Introduction
By contrast, here we focus on the development of a statistical mechanics approach that takes into account both complexity and functionality. Ultimately it is the circuit functionality - the specific computation implemented by a given circuit - that is the central object of interest in most computational problems. Our framework is based on reversible computing, which can be implemented either as permutations $P$ acting on the space of $2^{n}$ strings of $n$ bits, or as unitary transformations $U$ acting on the $d^{n}$ dimensional Hilbert space of $n$ qudits with local Hilbert space dimension $d$. For concreteness, in the body of the paper we focus on universal circuits implementing classical permutations for which the counting is discrete. Since practical implementations of quantum computation also use discrete sets of universal quantum gates to approximate continuum unitaries, the extension to quantum circuits is natural and is presented in the Supplementary Information.
@@
The circuit-thermodynamics framework has a natural application to the cryptographic concept of program obfuscation - a central problem in cryptography - which, in conjunction with other cryptographic primitives, underpins a tremendously powerful set of cryptographic tools. Program obfuscation is the process of “randomizing” a program while preserving its functionality. Specifically, given a computer program $P$, the process of obfuscating $P$ generates another program $P^{\prime}$ that (a) has a comparable size and the same functionality as $P$; and (b) $P^{\prime}$ provides no information about $P$ to an adversary with polynomial resources, except for its functionality [6]. Within circuit thermodynamics, the equilibrium state defined by a microcanonical ensemble in which an exponentially large number of ${\cal N}$-gate circuits with the same functionality appear with equal probability naturally realizes the obfuscation of every circuit in the distribution.
@@ Results
where $\Omega(P,{\cal N})$ is the number of circuits realizing permutation $P$ with exactly ${\cal N}$ gates. The latter definition implies (a) the inequality ${\cal N}\geq{\cal K}(P)$, where ${\cal K}(P)$ is the circuit complexity, i.e., the minimum number of gates required to implement the permutation $P$; and (b) the sum-rule $\sum_{P}\;\Omega(P,{\cal N})=|G|^{\cal N}$, where $|G|$ denotes the cardinality of the gate set used in the implementation of $P$.
@@
A number of inequalities follow from the definition of the circuit entropy in Eq. (1). The simplest one,
@@
expresses the fact that there may be more ways of implementing the product $P_{1}P_{2}$ than simply sequentially implementing $P_{1}$ and then $P_{2}$. In parallel with the entropy inequality in Eq. (2), the circuit complexity satisfies the opposite inequality,
@@
Using the inequality in Eq. (2), we can immediately derive a lower bound on the entropy ${\cal S}(P,{\cal N})$ in terms of the circuit complexity ${\cal K}(P)$ of the permutation $P$:
@@
where ${1\!\!1}$ denotes the identity permutation, which has zero complexity, i.e., it can be expressed without using any gate. Eq. (4) can be replaced with a more useful set of bounds, namely:
$\left[{\cal N}-{\cal K}(P)\right]\;\log_{2}|G|^{1/2}+{\cal S}(P,{\cal K}(P))\leq{\cal S}(P,{\cal N})<{\cal N}\;\log_{2}|G|\;.$ (5)
@@
Eq. (5) connects the microcanonical ensemble entropy to the circuit complexity and provides the foundation to what we refer to as the thermodynamics of circuit complexity.  [^2] In analogy to statistical physics, by introducing weights for different gates one could have also constructed a canonical ensemble for circuits. The two terms on the left hand side represent two types of contributions to the circuit entropy: the first depends on the “free volume” ${\cal N}-{\cal K}(P)$, with the complexity ${\cal K}(P)$ acting as an “excluded volume” and depends on $P$ only through the complexity ${\cal K}(P)$; the second is independent of ${\cal N}$ and accounts for the number of different possible circuits implementing the permutation $P$ within the minimum possible size ${\cal K}(P)$. We also note that the linear ${\cal N}$ dependence of both lower and upper bounds of ${\cal S}(P,{\cal N})$ in Eq. (5) implies that ${\cal S}(P,{\cal N})$ is extensive in ${\cal N}$. Moreover, the extensivity of ${\cal S}(P,{\cal N})$ with ${\cal N}$ and Eq. (4) evaluated for ${\cal N}={\cal K}(P)$ imply that ${\cal S}(P,{\cal K}(P))$ is at most extensive in ${\cal K}(P)$.
We proceed by making the following three assumptions, the last of which was motivated by the above discussion: (i) up to subextensive corrections, ${\cal S}(P,{\cal K}(P))$ depends on $P$ only through its complexity ${\cal K}(P)$; (ii) the entropy ${\cal S}(P,{\cal N})$ also only depends on ${\cal N}$ and the complexity ${\cal K}(P)$, namely,
${\cal S}(P,{\cal N})\approx\bar{\cal S}({\cal K}(P),{\cal N})\;.$ (6)
and that (iii) $\bar{\cal S}({\cal K}(P),{\cal N})$ is extensive in both ${\cal N}$ and ${\cal K}(P)$. Moreover, we assume that the free volume, ${\cal N}-{\cal K}(P)$, is an extensive thermodynamic variable and thus, the thermodynamic ensemble of circuits contains an exponential number of ${\cal N}$-gate circuits implementing any permutation $P$, as indicated by the inequality (5).
To finalize the formulation of circuit thermodynamics, we need to make one more assumption that pertains to a new entropy function defined as: $\sigma({\cal K},{\cal N})=\log_{2}\omega({\cal K},{\cal N})$. By contrast to $\Omega({\cal K}(P),{\cal N})$, which counts the number of ${\cal N}$-gate circuits of fixed functionality with complexity ${\cal K}$, $\omega({\cal K},{\cal N})$ counts all ${\cal N}$-gate circuits of fixed complexity, ${\cal K}$:
@@
leading to
$\sigma({\cal K},{\cal N})$ $\approx\log_{2}\nu({\cal K})+\bar{\cal S}({\cal K},{\cal N})\ \;,$ (8)
where $\nu({\cal K})=\sum_{P}\delta_{{\cal K},{\cal K}(P)}$ is a “density of states” that counts the number of possible functionalities (i.e., permutations) implemented by circuits of a fixed complexity, ${\cal K}$.
The last of the four assumptions that completely define the circuit thermodynamic framework is that $\sigma({\cal K},{\cal N})$ is an extensive function of ${\cal N}$ and ${\cal K}$. Extensivity with ${\cal N}$ simply follows from the extensivity of $\bar{\cal S}({\cal K},{\cal N})$ and Eq. (8). Extensivity with ${\cal K}$ implies that the density of states scales exponentially in ${\cal K}$, i.e., $\log_{2}\nu({\cal K})=\alpha{\cal K}$, consistent with the counting estimates for the number of circuits with given complexity ${\cal K}$ when degeneracies are ignored [2]. In the Supplementary Information we present a model for the growth of complexity of a random circuit, which accounts for functionality degeneracy, and recovers results obtained from the thermodynamic treatment that relies on the extensivity of $\bar{\cal S}({\cal K},{\cal N})$ and $\sigma({\cal K},{\cal N})$ with ${\cal N}$ and ${\cal K}$.
The function $\omega({\cal K},{\cal N})=2^{\sigma({\cal K},{\cal N})}$ defines the circuit complexity weight distribution for ${\cal N}$-gate circuits, which peaks at the extremum of the function $\sigma({\cal K},{\cal N})$:
$\frac{\partial{\sigma({\cal K},{\cal N})}}{\partial{\cal K}}{\biggr{\rvert}}_{{\cal N}}=0=\alpha+\frac{\partial\bar{\cal S}({\cal K},{\cal N})}{\partial{\cal K}}{\biggr{\rvert}}_{{\cal N}}\;.$ (9)
Since the entropy decreases with complexity, we write ${\partial\bar{\cal S}({\cal K},{\cal N})}/{\partial{\cal K}}|_{{\cal N}}=-\beta(\kappa)$ where $\beta(\kappa)$ is a positive function of $\kappa={\cal K}/{\cal N}$. At the extremum $\kappa=\kappa^{*}$, $\alpha=\beta(\kappa^{*})=\beta$, and Eq. (8) assumes the form,
@@
It is gratifying that Eq. (10) can be recast in a form familiar from the thermodynamics of physical systems: if we interpret ${\cal K}$ as the negative of the energy, $\mathcal{E}=-{\cal K}$ (or equivalently, $\epsilon=\mathcal{E}/{\cal N}=-\kappa$) and $\beta=1/{T}$ as the inverse of the temperature $T$, Eq. (10) can be rewritten in terms of the equilibrium free energy $F(T,{\cal N})$,
@@
where the form of $\omega({\cal K},{\cal N})$ can be obtained by expanding $\sigma({\cal K},{\cal N})=\log_{2}\omega({\cal K},{\cal N})=\alpha{\cal K}+\bar{\cal S}({\cal K},{\cal N})$ to second order in $\Delta{\cal K}=({\cal K}-{\cal K}^{*})$, the departure from the solution of the extremum condition, $\beta(\kappa^{*})=\alpha$:
$\log_{2}\omega({\cal K},{\cal N})=\bar{\cal S}({\cal K}^{*},{\cal N})+\beta{\cal K}^{*}+\frac{1}{2}\;{\Delta{\cal K}}^{2}\;\frac{\partial^{2}\bar{\cal S}({\cal K},{\cal N})}{\partial{\cal K}^{2}}{\biggr{\rvert}}_{{\cal N},{\cal K}={\cal K}^{*}}+\cdots\ \;.$ (13)
Using the extensivity of the entropy in ${\cal K}$ and ${\cal N}$ we can write $\bar{\cal S}({\cal K}^{*},{\cal N})=-\beta{\cal K}^{*}-\beta\mu{\cal N}+\cdots$, where we have borrowed the statistical mechanics notation involving the equilibrium “chemical potential” $\mu$. At the extremum, this leads to the simplification of the first two terms in Eq. (13) to $\bar{\cal S}({\cal K}^{*},{\cal N})+\alpha{\cal K}^{*}=-\beta\mu{\cal N}$ from which it then follows that:
@@
Thus, the probability distribution in Eq. (12) is a Gaussian peaked at ${\bar{\cal K}}={\cal K}^{*}=\kappa^{*}{\cal N}$ linear in ${\cal N}$ with a width (root mean square deviation) $\Delta_{\rm rms}\propto\sqrt{{\cal N}T^{2}c_{{\cal N}}}$, where $c_{{\cal N}}=-{\partial{\cal K}}/{\partial T}{|}_{{\cal K}^{*},{\cal N}}$ is a positive intensive quantity analogous to the specific heat in thermodynamics, which, in physical systems measures the increase in energy induced by an increase in temperature and controls energy fluctuations around the thermal equilibrium state at temperature $T$. In our case, note the negative sign in the definition of $c_{{\cal N}}$: this accounts for the fact that increasing temperature (i.e., increasing entropy) implies a decrease in complexity. Moreover, Eq. (33) and the sum rule $\sum_{P}\;\Omega(P,{\cal N})=\sum_{{\cal K}=0}^{{\cal N}}\omega({\cal K},{\cal N})=|G|^{\cal N}$ determines the leading behavior of the chemical potential, $\mu=-T\log_{2}|G|$.
@@
To further motivate the notion that generic circuits display a finite compressibility with a compression factor $10<\eta<$ 1, we consider the following scenario leading to a lower bound on $\eta$. Consider a random circuit of 3-bit Toffoli gates and imagine “pushing” a gate through the circuit until the gate encounters either (i) a gate with which it does not commute, in which case we stop; or (ii) its inverse, in which case the pair (the gate and its inverse) annihilate, decreasing the size of the circuit by two gates (see Fig. 1). For a Toffoli gate, the probability that it does commute with a gate on its path is of the order $1-{\cal O}(1/n)$, or equivalently, a gate can be “pushed” through ${\cal O}(n)$ gates before either stopping as in case (i) or annihilating as in case (ii). The overall probability of annihilation is ${\cal O}(1/n^{2})$, accounting for the probability ${\cal O}(1/n^{3})$ that the inverse is met in any of the ${\cal O}(n)$ attempts before stopping. Hence, this process leads to a compression of the circuit by a factor $(1-2\xi/n^{2})$ of its original size, where $\xi$ is a constant of ${\cal O}(1)$. This implies that circuits with greater than ${\cal O}(n^{2})$ universal (Toffoli) gates are compressible with $\eta=2\xi/n^{2}$, setting a lower bound for compressibility of random circuits of universal gates. Indeed, since the probability of annihilation of linear gates - NOTs and CNOTs - scales more favorably as $1/n$ and $1/n^{2}$ respectively, circuits comprised of gates from the universal set of Toffolis, CNOTs, and NOTs are more compressible with a compression factor $\eta$ above the Toffoli bound. [It is worth mentioning that since any linear circuit can be implemented with ${\cal O}(n^{2})$ gates, purely linear circuits are highly compressible, with $\eta\sim(1-n^{2}/{\cal N})$.]
@@
Thermodynamic mixing of circuits - an application to circuit obfuscation: As we have learned from physical systems, thermodynamics provides a platform for harnessing heat into useful work or for analyzing thermodynamic processes that transfer heat/entropy or molecular species. We find it reassuring that, in the context of circuits, this simplest of thermodynamic concepts - the entropy of mixing - provides a potentially important application to circuit obfuscation.  [^4] In traditional thermodynamic systems, the entropy of mixing provides detailed information about inter- and intra-species interactions in mixtures of different molecules and is used to produce uniform mixtures of chemicals. As already alluded to in the introduction, obfuscation of circuits follows immediately from two tenets of circuit thermodynamics: (i) the fact that there are exponentially many comparable size circuits which implement the same functionality; and (ii) that in the fully mixed maximum entropy state for an ensemble of such circuits is described by a uniform (microcanonical) distribution, i.e., drawing a circuit from this distribution results in any one of the exponentially many $\Omega(P,{\cal N})$ realizations of the ${\cal N}$-gate circuit with functionality $P$ with equal probability. Thus, any two different circuits of the same size and functionality drawn from the microcanonical ensemble of circuits cannot be distinguished from one another by an adversary with polynomial resources - the defining condition for the concept of circuit obfuscation.
@@
$T_{\left(g^{\,}_{i},\dots,g^{\,}_{i+k-1}\right),\left(g^{\prime}_{i},\dots,g^{\prime}_{i+k-1}\right)}=\frac{1}{\Omega(g_{i}\,\dots\,g_{i+k-1},k)}\;\delta_{g_{i}\,\dots\,g_{i+k-1},g^{\prime}_{i}\,\dots\,g^{\prime}_{i+k-1}}\;.$ (16)

### v3

@@
Abstract
@@
Circuit complexity, defined as the minimum circuit size required for implementing a particular Boolean computation, is a foundational concept in computer science. Determining circuit complexity is believed to be a hard computational problem [1, 2, 3]. Recently, in the context of black holes, circuit complexity has been promoted to a physical property, wherein the growth of complexity is reflected in the time evolution of the Einstein-Rosen bridge (“wormhole”) connecting the two sides of an AdS “eternal” black hole [4]. Here we are motivated by an independent set of considerations and explore links between complexity and thermodynamics for functionally-equivalent circuits, making the physics-inspired approach relevant to real computational problems, for which functionality is the key element of interest. In particular, our thermodynamic framework provides a new perspective on the obfuscation of programs of arbitrary length – an important problem in cryptography – as thermalization through recursive mixing of neighboring sections of a circuit, which can be viewed as the mixing of two containers with “gases of gates”. This recursive process equilibrates the average complexity and leads to the saturation of the circuit entropy, while preserving functionality of the overall circuit. The thermodynamic arguments hinge on ergodicity in the space of circuits which we conjecture is limited to disconnected ergodic sectors due to fragmentation. The notion of fragmentation has important implications for the problem of circuit obfuscation as it implies that there are circuits of same size and functionality that cannot be connected via a polynomial number of local moves. Furthermore, we argue that fragmentation is unavoidable unless the complexity classes NP and coNP coincide, a statement that implies the collapse of the polynomial hierarchy of computational complexity theory to its first level.
Here we propose a physics-inspired statistical mechanics approach to the behavior of reversible computational circuits of a given functionality, in which circuit complexity - the minimal number of gates required to implement that functionality - is treated as a thermodynamic variable. While over the past decade the connection between computational complexity and thermodynamics has emerged in the course of explorations of the quantum mechanics of black-holes [4, 5, 6], our own interest in this connection was motivated by a completely different and independent line of enquiry, namely one aimed at employing statistical mechanics techniques used to diagnose irreversibility and chaos in complex physical systems in the design of cryptographic tools for protecting confidential “data-in-use” – data and any manipulations of this data – from bounded adversaries [7, 8].
@@
By contrast to the discussion of black holes that focuses on the growth of complexity for very large random circuits and generally ignores the contribution of distinct circuits implementing the same functionality, the statistical mechanics approach presented here treats both complexity and functionality. Ultimately it is the circuit functionality - the specific computation implemented by a given circuit - that is the central object of interest in most computational problems. Our framework is based on reversible computing, which can be implemented either as permutations $P$ acting on the space of $2^{n}$ strings of $n$ bits, or as unitary transformations $U$ acting on the $d^{n}$ dimensional Hilbert space of $n$ qudits with local Hilbert space dimension $d$. For concreteness, in the body of the paper we focus on universal circuits implementing classical permutations for which the counting is discrete.  [^1] We note that here we focus on even permutations in the alternating group $A_{2^{n}}$, which can be decomposed into products of universal 3-bit gates [9]. The realization of odd permutations requires either one additional $n$-bit gate or one ancilla bit. More generally, the group $S_{2^{n}}$ can be embedded in the alternating group $A_{2^{n+1}}$, an embedding that formalizes the incorporation of ancillas, and which also suggests that, given a computational model, the thermodynamic discussion presented here for even permutation extends to odd permutations without essential changes. Since practical implementations of quantum computation also use discrete sets of universal quantum gates to approximate continuum unitaries, the extension to quantum circuits is natural and is presented in Section A of the Supplementary Information.
@@
The circuit-thermodynamics framework has a natural application to the cryptographic concept of program obfuscation - a central problem in cryptography - which, in conjunction with other cryptographic primitives, enables a tremendously powerful set of cryptographic tools. Program obfuscation is the process of “randomizing” a program while preserving its functionality. Specifically, given a computer program $P$, the process of obfuscating $P$ generates another program $P^{\prime}$ that (a) has a comparable size and the same functionality as $P$; and (b) $P^{\prime}$ provides no information about $P$ to an adversary with polynomial resources, except for its functionality [10]. Within circuit thermodynamics, the equilibrium state defined by a microcanonical ensemble in which an exponentially large number of ${\cal N}$-gate circuits with the same functionality appear with equal probability naturally realizes the obfuscation of every circuit in the distribution.
@@ Results
where $\Omega(P,{\cal N})$ is the number of circuits realizing a permutation $P$ with exactly ${\cal N}$ gates. The latter definition implies (a) the inequality ${\cal N}\geq{\cal K}(P)$, where ${\cal K}(P)$ is the circuit complexity, i.e., the minimum number of gates required to implement the permutation $P$; and (b) the sum-rule $\sum_{P}\;\Omega(P,{\cal N})=|G|^{\cal N}$, where $|G|$ denotes the cardinality of the gate set used in the implementation of $P$.  [^2] For 3-bit gates in $S_{8}$, $|G|=8!\binom{n}{3}$ and for gates in the universal set (NOT, CNOT, Toffoli), $|G|=n+n(n-1)+n(n-1)(n-2)/2$. We expect that, as is the case in physical systems, our circuit thermodynamic description is universal, with details of the gate set only appearing in the values of (intensive) non-universal parameters.
@@
A number of inequalities follow from the definition of the circuit entropy in  (1). The simplest one,
@@
expresses the fact that there may be more ways of implementing the product $P_{1}P_{2}$ than simply sequentially implementing $P_{1}$ and then $P_{2}$. In parallel with the entropy inequality in  (2), the circuit complexity satisfies the opposite inequality,
@@
Using the inequality in  (2), we can immediately derive a lower bound on the entropy ${\cal S}(P,{\cal N})$ in terms of the circuit complexity ${\cal K}(P)$ of the permutation $P$:
@@
where ${1\!\!1}$ denotes the identity permutation, which has zero complexity, i.e., it can be expressed without using any gate.  (4) can be replaced with a more useful set of bounds, namely:
$\left[{\cal N}-{\cal K}(P)\right]\;\log_{2}|G|^{1/2}+{\cal S}(P,{\cal K}(P))$ $\leq{\cal S}(P,{\cal N})$
$<{\cal N}\;\log_{2}|G|\;.$ (5)
@@
Finally, the last upper bound in  (5) follows immediately from the sum rule $\sum_{P}\;\Omega(P,{\cal N})=|G|^{\cal N}$, which implies $\Omega(P,{\cal N})<|G|^{\cal N}$ and ${\cal S}(P,{\cal N})<{\cal N}\;\log_{2}|G|$.
Equation (5) and the discussion accompanying it connect the microcanonical ensemble entropy to the circuit complexity and provide the foundation to what we refer to as the thermodynamics of circuit complexity.  [^3] In analogy to statistical physics, by introducing weights for different gates one could have also constructed a canonical ensemble for circuits. In particular, the linear ${\cal N}$ dependence of both lower and upper bounds of ${\cal S}(P,{\cal N})$ in  (5) implies that ${\cal S}(P,{\cal N})$ is extensive in ${\cal N}$. Moreover, the assumed extensivity in ${\cal K}(P)$ of the lower bound on ${\cal S}(P,{\cal K}(P))$ motivated in the Supplementary Information, together with the upper bound ${\cal S}(P,{\cal K}(P))<(\log_{2}|G|){\cal K}(P)$ obtained from (5) evaluated for ${\cal N}={\cal K}(P)$ imply that ${\cal S}(P,{\cal K}(P))$ is extensive in ${\cal K}(P)$, a property we express as
${\cal S}(P,{\cal K}(P)\approx\gamma(P;G)\;{\cal K}(P)\;,$ (6)
where $\gamma(P;G)$ depends on the permutation $P$ and the gate set, $G$, used in implementing the permutation $P$.
The main assumption of our approach, motivated by the above discussion, is that ${\cal S}(P,{\cal N})$ is an extensive function of both ${\cal N}$ and ${\cal K}(P)$ an assumption analogous to that used in the microcanonical ensemble derivation of thermodynamics for physical systems in which the extensive thermodynamic quantities are the number of particles and the energy. An important consequence of the extensivity of the entropy with the “free volume” ${\cal N}-{\cal K}$ is that the thermodynamic ensemble of ${\cal N}$-gate circuits contains a number of circuits $\Omega(P,{\cal N})$ implementing any permutation $P$ that scales exponentially with the ${\cal N}-{\cal K}(P)$, behavior we will use repeatedly in what follows.
To finalize the formulation of circuit thermodynamics, we introduce the function $\omega({\cal K},{\cal N})$ which, by contrast to $\Omega(P,{\cal N})$, which counts the number of ${\cal N}$-gate circuits of fixed functionality, counts all ${\cal N}$-gate circuits of fixed complexity, ${\cal K}$:
$\omega({\cal K},{\cal N})$ $=\sum_{P}\delta_{{\cal K},{\cal K}(P)}\;\Omega(P,{\cal N})$
@@
[…]
where $\nu({\cal K})=\sum_{P}\delta_{{\cal K},{\cal K}(P)}$ is a “density of states” that counts the number of possible functionalities (i.e., permutations) implemented by circuits of a fixed complexity, ${\cal K}$; and $\bar{\cal S}({\cal K},{\cal N})$ defines an “annealed average” of ${\cal S}({\cal K},{\cal N})$ over $P$. In turn, this allows us to write:
$\sigma({\cal K},{\cal N})=\log_{2}\omega({\cal K},{\cal N})$ $\equiv\log_{2}\nu({\cal K})+\bar{\cal S}({\cal K},{\cal N})\ \;.$ (8)
The extensively of ${\cal S}(P,{\cal N})$ with ${\cal N}$ and ${\cal K}(P)$ implies that $\bar{\cal S}({\cal K},{\cal N})$ is itself extensive in ${\cal N}$ and ${\cal K}$. The function $\omega({\cal K},{\cal N})=2^{\sigma({\cal K},{\cal N})}$, which defines the circuit complexity weight distribution for ${\cal N}$-gate circuits, peaks at the extremum of the function $\sigma({\cal K},{\cal N})$:
$\frac{\partial{\sigma({\cal K},{\cal N})}}{\partial{\cal K}}{\biggr{\rvert}}_{{\cal N}}=0=\frac{\partial{\log_{2}\nu({\cal K})}}{\partial{\cal K}}+\frac{\partial\bar{\cal S}({\cal K},{\cal N})}{\partial{\cal K}}{\biggr{\rvert}}_{{\cal N}}\;.$ (9)
Since the entropy $\bar{\cal S}({\cal K},{\cal N})$ is an extensive decreasing function of complexity, ${\partial\bar{\cal S}({\cal K},{\cal N})}/{\partial{\cal K}}|_{{\cal N}}=-\beta(\kappa)$, where $\beta(\kappa)$ is a positive intensive function of $\kappa={\cal K}/{\cal N}$. At the extremum, $\kappa=\kappa^{*}$, $\beta\equiv\beta(\kappa^{*})$, and thus ${\partial\;{\log_{2}\nu({\cal K})}}/{\partial{\cal K}}|_{{\cal K}^{*}}=\beta$, implying that $\nu({\cal K}^{*})=2^{\beta{\cal K}^{*}}$ increases exponentially with complexity, consistent with the counting estimates for the number of circuits with given complexity ${\cal K}$ [4]. At the extremum,  (8) assumes the form
@@
It is gratifying that  (10) can be recast in a form familiar from the thermodynamics of physical systems: if we interpret ${\cal K}$ as the negative of the energy, $\mathcal{E}=-{\cal K}$ (or equivalently, $\epsilon=\mathcal{E}/{\cal N}=-\kappa$) and $\beta=1/{T}$ as the inverse of the temperature $T$,  (10) can be rewritten in terms of the equilibrium free energy $F(T,{\cal N})$,
@@
where the form of $\omega({\cal K},{\cal N})$ can be obtained by expanding $\sigma({\cal K},{\cal N})=\log_{2}\omega({\cal K},{\cal N})=\beta{\cal K}+\bar{\cal S}({\cal K},{\cal N})$ to second order in $\Delta{\cal K}=({\cal K}-{\cal K}^{*})$, the departure from the solution of the extremum condition:
$\log_{2}\omega({\cal K},{\cal N})$ $=\bar{\cal S}({\cal K}^{*},{\cal N})+\beta{\cal K}^{*}$
$\;\;+\frac{1}{2}\;{\Delta{\cal K}}^{2}\;\frac{\partial^{2}\bar{\cal S}({\cal K},{\cal N})}{\partial{\cal K}^{2}}{\biggr{\rvert}}_{{\cal N},{\cal K}={\cal K}^{*}}+\cdots\ \;.$ (13)
Using the extensivity of the entropy in ${\cal K}$ and ${\cal N}$ we can write $\bar{\cal S}({\cal K}^{*},{\cal N})=-\beta{\cal K}^{*}-\beta\mu{\cal N}+\cdots$, where we have borrowed the statistical mechanics notation involving the equilibrium “chemical potential” $\mu$. At the extremum, this leads to the simplification of the first two terms in  (13) to $\bar{\cal S}({\cal K}^{*},{\cal N})+\beta{\cal K}^{*}=-\beta\mu{\cal N}$ from which it then follows that:
@@
Thus, the probability distribution in  (12) is a Gaussian peaked at ${\cal K}^{*}=\kappa^{*}{\cal N}$ linear in ${\cal N}$ with a width (root mean square deviation) $\Delta_{\rm rms}\propto\sqrt{{\cal N}T^{2}c_{{\cal N}}}$, where $c_{{\cal N}}=-{\partial{\cal K}}/{\partial T}{|}_{{\cal K}^{*},{\cal N}}$ is a positive intensive quantity analogous to the specific heat in thermodynamics, which, in physical systems measures the increase in energy induced by an increase in temperature and controls energy fluctuations around the thermal equilibrium state at temperature $T$. Finally,  (14) and the sum rule $\sum_{P}\;\Omega(P,{\cal N})=\sum_{{\cal K}=0}^{{\cal N}}\omega({\cal K},{\cal N})=|G|^{\cal N}$ determines the leading behavior of the chemical potential, $\mu=-T\log_{2}|G|$.
@@
To further motivate the notion that generic circuits display a finite compressibility with a compression factor $10<\eta<$ 1, we consider the following scenario leading to a lower bound on $\eta$. Consider a random circuit of 3-bit Toffoli gates and imagine “pushing” a gate through the circuit until the gate encounters either (i) a gate with which it does not commute, in which case we stop; or (ii) its inverse, in which case the pair (the gate and its inverse) annihilate, decreasing the size of the circuit by two gates (see Fig. 1). [^5] Note that here we are considering random circuits built gate-by-gate, as opposed to minimum circuits describing random permutations. The former are compressible while the latter are not. For a Toffoli gate, the probability that it does commute with a gate on its path is of the order $1-{\cal O}(1/n)$, or equivalently, a gate can be “pushed” through ${\cal O}(n)$ gates before either stopping as in case (i) or annihilating as in case (ii). The overall probability of annihilation is ${\cal O}(1/n^{2})$, accounting for the probability ${\cal O}(1/n^{3})$ that the inverse is met in any of the ${\cal O}(n)$ attempts before stopping. Hence, this process leads to a compression of the circuit by a factor $(1-2\xi/n^{2})$ of its original size, where $\xi$ is a constant of ${\cal O}(1)$. This implies that circuits with greater than ${\cal O}(n^{2})$ universal (Toffoli) gates are compressible with $\eta=2\xi/n^{2}$, setting a lower bound for compressibility of random circuits of universal gates. Indeed, since the probability of annihilation of linear gates - NOTs and CNOTs - scales more favorably as $1/n$ and $1/n^{2}$ respectively, circuits comprised of gates from the universal set of Toffolis, CNOTs, and NOTs are more compressible with a compression factor $\eta$ above the Toffoli bound. [It is worth mentioning that since any linear circuit can be implemented with ${\cal O}(n^{2})$ gates, purely linear circuits are highly compressible, with $\eta\sim(1-n^{2}/{\cal N})$.]
@@
Thermodynamic mixing of circuits - an application to circuit obfuscation: As we have learned from physical systems, thermodynamics provides a platform for harnessing heat into useful work or for analyzing thermodynamic processes that transfer heat/entropy or molecular species. We find it reassuring that, in the context of circuits, this simplest of thermodynamic concepts - the entropy of mixing - provides a potentially important application to circuit obfuscation. As already alluded to in the introduction, obfuscation of circuits follows immediately from two tenets of circuit thermodynamics: (i) the fact that there are exponentially many comparable size circuits which implement the same functionality; and (ii) that in the fully mixed maximum entropy state for an ensemble of such circuits is described by a uniform (microcanonical) distribution, i.e., drawing a circuit from this distribution results in any one of the exponentially many $\Omega(P,{\cal N})$ realizations of the ${\cal N}$-gate circuit with functionality $P$ with equal probability. Thus, any two different circuits of the same size and functionality drawn from the microcanonical ensemble of circuits cannot be distinguished from one another by an adversary with polynomial resources - the defining condition for the concept of circuit obfuscation.
@@
$T_{\left(g^{\,}_{i},\dots,g^{\,}_{i+k-1}\right),\left(g^{\prime}_{i},\dots,g^{\prime}_{i+k-1}\right)}$
$\hskip 28.45274pt=\frac{1}{\Omega(g_{i}\,\dots\,g_{i+k-1},k)}\;\delta_{g_{i}\,\dots\,g_{i+k-1},g^{\prime}_{i}\,\dots\,g^{\prime}_{i+k-1}}\;.$ (16)